400 tickets were sold to a ballgame. Adult tickets were $2, Child tickets were $3. Tickets sold totaled $1050. How many of each ticket were sold.
Answer by kunal_122(19)

2x+(400-x)*3=1050
2x+1200-3x=1050
-x=-150
x=150
400-x=250
just substitue these number and u will get 1050=1050 which means that the answer is correct